OpenStudy (anonymous):

When the positive integer x is divided by 5, the remainder is 3. What is the remainder when x+19 is divided by 5?

OpenStudy (anonymous):

where did you get that equation from?

OpenStudy (lxelle):

I'm using the division of polynomial form P(x) = D(x)Q(x) + R(x)

OpenStudy (anonymous):

oh ok, i've never heard of that so i was confused

OpenStudy (anonymous):

i'm in Pre Calculas in High School (11th grade)

OpenStudy (irishboy123):

x = 5n + 3 (n = integer) (x+19)/5 = (5n+3+19)/5 = (5n+22)/5 = n + 4 rem 2